# Benchmarking and Improving Detail Image Caption

### v3

@@ Abstract
Image captioning has long been regarded as a fundamental task in visual understanding. Recently, however, few large vision-language model (LVLM) research discusses model’s image captioning performance because of the outdated short-caption benchmarks and unreliable evaluation metrics. In this work, we propose to benchmark detail image caption task by curating high-quality evaluation datasets annotated by human experts, GPT-4V and Gemini-1.5-Pro. We also design a more reliable caption evaluation metric called CAPTURE (CAPtion evaluation by exTracting and coUpling coRE information). CAPTURE extracts visual elements, e.g., objects, attributes and relations from captions, and then matches these elements through three stages, achieving the highest consistency with expert judgements over other rule-based or model-based caption metrics. The proposed benchmark and metric provide reliable evaluation for LVLM’s detailed image captioning ability. Guided by this evaluation, we further explore to unleash LVLM’s detail caption capabilities by synthesizing high-quality data through a five-stage data construction pipeline. Our pipeline only uses a given LVLM itself and other open-source tools, without any human or GPT-4V annotation in the loop. Experiments show that the proposed data construction strategy significantly improves model-generated detail caption data quality for LVLMs with leading performance, and the data quality can be further improved in a self-looping paradigm. All code and dataset will be publicly available at https://github.com/foundation-multimodal-models/CAPTURE.
@@ 1 Introduction
Image captioning has long been a fundamental task to assess LVLM’s vision understanding capability [54, 34, 12, 15]. However, recent LVLM researches evaluate LVLMs’ visual understanding performance with a focus on QA benchmarks, such as MME [16], MMBench [36], MMMU [60], MM-Vet [59], etc., which may suffer from instability caused by LVLMs’ varying instruction following abilities [16]. What’s worse, human-defined queries may cover a limited scope of vision features [25] and introduce bias in performance evaluation [59]. Traditional image captioning task is considered unreliable for visual understanding evaluation because of the outdated benchmarks and unstable evaluation metrics. Current image caption benchmarks consist of fairly short captions with limited vision information [32, 2], while SOTA LVLMs are capable of generating detail image captions encompassing a variety of fine-grained elements [9, 54, 34], and only a few of them are covered in the provided ground truth captions. This contradiction leads to unsatisfying evaluation results. To this end, we propose to curate high-quality detail image caption evaluation datasets to provide reliable evaluation results for SOTA LVLMs. The evaluation datasets are annotated by human experts and the most capable LVLM GPT-4V[41]&Gemini-1.5-Pro[46], and are therefore of satisfying quality for state-of-the-art (SOTA) LVLM evaluation.
@@
(1) We release a 4870-case GPT-4V&Gemini-1.5-Pro annotated detail image caption benchmark for reliable evaluation, accompanied with three model-generated captions and corresponding GPT-4 annotated quality scores for expert judgement consistency evaluation.
(2) We propose a novel detail image caption evaluation metric CAPTURE, which adopts a T5-based parser to extract visual elements from captions, and compute the matching score via a three-stage matching module. Experiments indicate that CAPTURE metric achieves the highest consistency with human and GPT-4V&Gemini-1.5-Pro judgement over other caption metrics, providing reliable detail caption evaluation results without expensive LLM API calls.
@@ 3.1 Detail Caption Evaluation Datasets
DetailCaps-4870 COCO [32], SAM [23], LAION [49] GPT-4V $4870$ $9740$ $122.06$ $377,184$
@@
We further curate a 4870 case dataset annotated by GPT-4V&Gemini-1.5-Pro for detail caption evaluation. Besides the data sources used in human-annotated 100 cases, we further incorporate pictures from COYO [6], LAION [49], CC [7] and Flickr [58] for diversity. Captions generated by ShareCaptioner, CogVLM and LLaVA-1.5 and corresponding annotated caption scores are provided for each sample. We instruct text-only GPT-4[1] to compare model-generated captions with GPT-4V&Gemini-1.5-Pro annotated references to obtain evaluation scores. We use text-only GPT-4 for evaluation because of its outstanding instruction following abilities. We refer to Appendix A for more details about the prompts used for detail caption generation and GPT4 evaluation generation.
@@ 5.1.2 Main Results
BLEU [2002] $0.2150$ $96.27$ $0.1623$ $0.2163$ $0.3099$ $24.86$ $0.2135$ $0.2812$ $0.2625$ $60.57$ $0.1879$ $0.2488$
ROUGE [2004] $0.2554$ $185.69$ $0.1905$ $0.3321$ $0.3291$ $90.42$ $0.2348$ $0.3303$ $0.2923$ $138.06$ $0.2127$ $0.3312$
METEOR [2005] $0.3643$ $384.58$ $0.2679$ $0.3529$ $0.4386$ $193.46$ $0.3165$ $0.4621$ $0.4015$ $289.02$ $0.2922$ $0.4075$
CIDER [2015] $0.0834$ $1.7e^{7}$ $0.1159$ $0.0564$ $0.1213$ $2.27e^{7}$ $0.0908$ $0.0948$ $0.1024$ $1.99e^{7}$ $0.1034$ $0.0756$
@@
SPICE [2016] $0.3580$ $126.60$ $0.2641$ $0.3819$ $0.5155$ $131.1$ $0.3818$ $0.5554$ $0.4368$ $128.85$ $0.3230$ $0.4687$
CLIP-Score [2021] $0.2532$ $48.81$ $0.1807$ $0.2928$ $0.4463$ $16.10$ $0.3039$ $0.4109$ $0.3498$ $32.46$ $0.2423$ $0.3519$
PAC-S [2023] $0.2584$ $62.82$ $0.1833$ $0.2843$ $0.2783$ $18.93$ $0.1795$ $0.2930$ $0.2684$ $40.88$ $0.1814$ $0.2887$
CAPTURE $\bm{0.4735}$ $\bm{11.58}$ $\bm{0.3688}$ $\bm{0.6117}$ $\bm{0.5366}$ $\bm{4.82}$ $\bm{0.3956}$ $\bm{0.5737}$ $\bm{0.5051}$ $\bm{8.20}$ $\bm{0.3822}$ $\bm{0.5927}$
@@ 5.1.3 Analysis
CAPTURE $0.4735$ $11.58$ $0.3688$ $0.6117$ $0.5366$ $4.82$ $0.3956$ $0.5737$ $0.5051$ $8.20$ $0.3822$ $0.5927$
   - Stop words $0.4830$ $13.23$ $0.3804$ $0.5947$ $0.5356$ $5.94$ $0.3940$ $0.5730$ $0.5093$ $9.58$ $0.3872$ $0.5838$
   - Soft matching $0.4674$ $29.15$ $0.3488$ $0.5770$ $0.5426$ $18.65$ $0.3964$ $0.5685$ $0.5050$ $23.90$ $0.3726$ $0.5728$
   $\alpha,\beta,\gamma=5,5,0$ $0.4654$ $9.21$ $0.3642$ $0.5947$ $0.5282$ $3.94$ $0.3941$ $0.5686$ $0.4968$ $6.58$ $0.3792$ $0.5816$
   $\alpha,\beta,\gamma=5,5,5$ $0.4651$ $13.75$ $0.3556$ $0.6064$ $0.5270$ $5.65$ $0.3834$ $0.5659$ $0.4960$ $9.70$ $0.3695$ $0.5862$
@@
CogVLM[2023a] Vicuna-7B Human Annt. $490^{2}$ $63.01$ $60.20$
ShareCaptioner-7B[2023a] Vicuna-7B GPT-4V Annt. $448^{2}$ $60.85$ $59.55$
LLaVA-1.5-7B[2023a] Vicuna-7B Synthesized $336^{2}$ $51.23$ $49.69$
LLaVA-1.5-13B[2023a] Vicuna-13B Synthesized $336^{2}$ $51.74$ $51.10$
LLaVA-NEXT-7B[2024a] Vicuna-7B GPT-4V Annt. $336^{2}$*{$1$-$5$} $60.18$ $58.22$
LLaVA-NEXT-13B[2024a] Vicuna-13B GPT-4V Annt. $336^{2}$*{$1$-$5$} $60.38$ $58.66$
LLaVA-NEXT-34B[2024a] Hermes-2-Yi-34B GPT-4V Annt. $336^{2}$*{$1$-$5$} $60.60$ $58.88$
Mini-Gemini-HD-7B[2024] Vicuna-7B GPT-4V Annt. $336^{2}$*$5$ $59.51$ $57.58$
Mini-Gemini-HD-13B[2024] Vicuna-13B GPT-4V Annt. $336^{2}$*$5$ $60.51$ $58.39$
Intern-XComposerV2[2024] Vicuna-7B GPT-4V Annt. $490^{2}$ $61.43$ $59.92$
InternVL-V1.2-PLUS-40B[2023b] Hermes-2-Yi-34B GPT-4V Annt. $448^{2}$ $61.61$ $60.75$
InternVL-V1.5-26B[2024c] InternLM-20B GPT-4V Annt. $448^{2}$*{$1$-$41$} $65.62$ $62.36$
@@ 5.2.3 Analysis
Self $51.23$ $65.24$ $43.31$ $51.27$ $64.61$ $43.79$ $51.25$ $64.92$ $43.55$
Synthesized $57.11$ $66.31$ $52.16$ $56.18$ $63.02$ $51.48$ $56.64$ $64.67$ $51.82$
   - filter $56.78$ $65.16$ $53.26$ $56.01$ $62.80$ $51.51$ $56.39$ $63.98$ $52.38$
   vqa filter $56.44$ $63.95$ $51.11$ $55.85$ $62.80$ $51.12$ $56.14$ $63.38$ $51.11$
   filter local $56.75$ $63.87$ $51.61$ $56.24$ $62.74$ $51.90$ $56.50$ $63.30$ $51.75$
@@
Self $51.23$ $65.24$ $43.31$ $51.27$ $64.61$ $43.79$ $51.25$ $64.92$ $43.55$
   loop1 $51.91$ $63.48$ $45.02$ $52.52$ $63.76$ $45.81$ $52.22$ $63.62$ $45.42$
   loop2 $52.50$ $63.43$ $45.66$ $52.53$ $62.57$ $46.42$ $52.52$ $63.00$ $46.04$
   loop3 $52.89$ $62.45$ $46.86$ $52.86$ $61.83$ $47.25$ $52.88$ $62.14$ $47.05$
   loop4 $54.02$ $62.24$ $48.45$ $54.38$ $61.68$ $49.50$ $54.20$ $61.96$ $48.98$

### v4

@@ Abstract
Image captioning has long been regarded as a fundamental task in visual understanding. Recently, however, few large vision-language model (LVLM) research discusses model’s image captioning performance because of the outdated short-caption benchmarks and unreliable evaluation metrics. In this work, we propose to benchmark detail image caption task by curating high-quality evaluation datasets annotated by human experts, GPT-4V, Gemini-1.5-Pro and GPT-4O. We also design a more reliable caption evaluation metric called CAPTURE (CAPtion evaluation by exTracting and coUpling coRE information). CAPTURE extracts visual elements, e.g., objects, attributes and relations from captions, and then matches these elements through three stages, achieving the highest consistency with expert judgements over other rule-based or model-based caption metrics. The proposed benchmark and metric provide reliable evaluation for LVLM’s detailed image captioning ability. Guided by this evaluation, we further explore to unleash LVLM’s detail caption capabilities by synthesizing high-quality data through a five-stage data construction pipeline. Our pipeline only uses a given LVLM itself and other open-source tools, without any human or GPT-4V annotation in the loop. Experiments show that the proposed data construction strategy significantly improves model-generated detail caption data quality for LVLMs with leading performance, and the data quality can be further improved in a self-looping paradigm. All code and dataset will be publicly available at https://github.com/foundation-multimodal-models/CAPTURE.
@@ 1 Introduction
Image captioning has long been a fundamental task to assess LVLM’s vision understanding capability [55, 34, 12, 16]. However, recent LVLM researches evaluate LVLMs’ visual understanding performance with a focus on QA benchmarks, such as MME [17], MMBench [36], MMMU [61], MM-Vet [60], etc., which may suffer from instability caused by LVLMs’ varying instruction following abilities [17]. What’s worse, human-defined queries may cover a limited scope of vision features [25] and introduce bias in performance evaluation [60]. Traditional image captioning task is considered unreliable for visual understanding evaluation because of the outdated benchmarks and unstable evaluation metrics. Current image caption benchmarks consist of fairly short captions with limited vision information [32, 2], while SOTA LVLMs are capable of generating detail image captions encompassing a variety of fine-grained elements [9, 55, 34], and only a few of them are covered in the provided ground truth captions. This contradiction leads to unsatisfying evaluation results. To this end, we propose to curate high-quality detail image caption evaluation datasets to provide reliable evaluation results for SOTA LVLMs. The evaluation datasets are annotated by human experts and the most capable LVLM GPT-4V[41], Gemini-1.5-Pro[47] and GPT-4O [42], and are therefore of satisfying quality for state-of-the-art (SOTA) LVLM evaluation.
@@
(1) We release a 4870-case GPT-4V, Gemini-1.5-Pro and GPT-4O annotated detail image caption benchmark for reliable evaluation, accompanied with three model-generated captions and corresponding GPT-4 annotated quality scores for expert judgement consistency evaluation.
(2) We propose a novel detail image caption evaluation metric CAPTURE, which adopts a T5-based parser to extract visual elements from captions, and compute the matching score via a three-stage matching module. Experiments indicate that CAPTURE metric achieves the highest consistency with expert judgement over other caption metrics, providing reliable detail caption evaluation results without expensive LLM API calls.
@@ 3.1 Detail Caption Evaluation Datasets
DetailCaps-4870 COCO [32], SAM [23], LAION [50] GPT-4V, GPT-4O $4870$ $14610$ $122.06$ $533,201$
@@
We further curate a 4870 case dataset annotated by GPT-4V, Gemini-1.5-Pro and GPT-4O for detail caption evaluation. Besides the data sources used in human-annotated 100 cases, we further incorporate pictures from COYO [6], LAION [50], CC [7] and Flickr [59] for diversity. Captions generated by ShareCaptioner, CogVLM and LLaVA-1.5 and corresponding annotated caption scores are provided for each sample. We instruct text-only GPT-4[1] to compare model-generated captions with GPT-4V, Gemini-1.5-Pro and GPT-4O annotated references to obtain evaluation scores. We use text-only GPT-4 for evaluation because of its outstanding instruction following abilities. We refer to Appendix A for more details about the prompts used for detail caption generation and GPT4 evaluation generation.
@@ 5.1.2 Main Results
BLEU [2002] $0.2150$ $96.27$ $0.1623$ $0.2163$ $0.3066$ $13.24$ $0.2109$ $0.2760$ $0.2608$ $54.75$ $0.1866$ $0.2462$
ROUGE [2004] $0.2554$ $185.69$ $0.1905$ $0.3321$ $0.3347$ $82.55$ $0.2393$ $0.3445$ $0.2951$ $134.12$ $0.2149$ $0.3383$
METEOR [2005] $0.3643$ $384.58$ $0.2679$ $0.3529$ $0.4400$ $196.19$ $0.3175$ $0.4594$ $0.4022$ $290.38$ $0.2927$ $0.4062$
CIDER [2015] $0.0834$ $1.7e^{7}$ $0.1159$ $0.0564$ $0.1462$ $3.5e^{7}$ $0.1171$ $0.1418$ $0.1148$ $2.6e^{7}$ $0.1165$ $0.0991$
@@
SPICE [2016] $0.3580$ $126.60$ $0.2641$ $0.3819$ $0.5192$ $185.30$ $0.3847$ $0.5616$ $0.4386$ $155.95$ $0.3244$ $0.4718$
RefCLIPScore [2021] $0.2538$ $31.82$ $0.1829$ $0.3244$ $0.4577$ $11.11$ $0.3129$ $0.4437$ $0.3558$ $21.46$ $0.2479$ $0.3841$
RefPAC-S [2023] $0.2664$ $60.67$ $0.1946$ $0.3221$ $0.4135$ $19.95$ $0.3246$ $0.3825$ $0.3399$ $40.31$ $0.2596$ $0.3523$
CAPTURE $\bm{0.4735}$ $\bm{11.58}$ $\bm{0.3688}$ $\bm{0.6117}$ $\bm{0.5446}$ $\bm{5.00}$ $\bm{0.4033}$ $\bm{0.5919}$ $\bm{0.5091}$ $\bm{8.29}$ $\bm{0.3861}$ $\bm{0.6018}$
@@ 5.1.3 Analysis
CAPTURE $0.4735$ $11.58$ $0.3688$ $0.6117$ $0.5446$ $5.00$ $0.4033$ $0.5919$ $0.5091$ $8.29$ $0.3861$ $0.6018$
   - Stop words $0.4830$ $13.23$ $0.3804$ $0.5947$ $0.5456$ $6.13$ $0.4047$ $0.5859$ $0.5143$ $9.68$ $0.3926$ $0.5903$
   - Soft matching $0.4674$ $29.15$ $0.3488$ $0.5770$ $0.5616$ $20.35$ $0.4116$ $0.5914$ $0.5145$ $24.75$ $0.3802$ $0.5842$
   $\alpha,\beta,\gamma=5,5,0$ $0.4654$ $9.21$ $0.3642$ $0.5947$ $0.5335$ $4.05$ $0.4002$ $0.5802$ $0.4994$ $6.63$ $0.3822$ $0.5875$
   $\alpha,\beta,\gamma=5,5,5$ $0.4651$ $13.75$ $0.3556$ $0.6064$ $0.5388$ $5.92$ $0.3936$ $0.5844$ $0.5020$ $9.84$ $0.3746$ $0.5954$
   $\alpha,\beta,\gamma=3,7,2$ $0.4842$ $10.59$ $0.3863$ $0.5654$ $0.5308$ $5.17$ $0.3993$ $0.5846$ $0.5075$ $7.88$ $0.3928$ $0.5750$
   $\alpha,\beta,\gamma=7,3,2$ $0.4384$ $10.10$ $0.3458$ $0.6010$ $0.5231$ $4.36$ $0.3874$ $0.5698$ $0.4808$ $7.23$ $0.3666$ $0.5854$
@@
CogVLM[2023a] Vicuna-7B Human Annt. $490^{2}$ $63.01$ $60.06$
ShareCaptioner-7B[2023a] Vicuna-7B GPT-4V Annt. $448^{2}$ $60.85$ $59.80$
LLaVA-1.5-7B[2023a] Vicuna-7B Synthesized $336^{2}$ $51.23$ $51.05$
LLaVA-1.5-13B[2023a] Vicuna-13B Synthesized $336^{2}$ $51.74$ $51.20$
LLaVA-NEXT-7B[2024a] Vicuna-7B GPT-4V Annt. $336^{2}$*{$1$-$5$} $60.18$ $58.61$
LLaVA-NEXT-13B[2024a] Vicuna-13B GPT-4V Annt. $336^{2}$*{$1$-$5$} $60.38$ $59.01$
LLaVA-NEXT-34B[2024a] Hermes-2-Yi-34B GPT-4V Annt. $336^{2}$*{$1$-$5$} $60.60$ $59.20$
Mini-Gemini-HD-7B[2024] Vicuna-7B GPT-4V Annt. $336^{2}$*$5$ $59.51$ $57.95$
Mini-Gemini-HD-13B[2024] Vicuna-13B GPT-4V Annt. $336^{2}$*$5$ $60.51$ $58.66$
Intern-XComposerV2[2024] Vicuna-7B GPT-4V Annt. $490^{2}$ $61.43$ $59.86$
InternVL-V1.2-PLUS-40B[2023b] Hermes-2-Yi-34B GPT-4V Annt. $448^{2}$ $61.61$ $60.69$
InternVL-V1.5-26B[2024c] InternLM-20B GPT-4V Annt. $448^{2}$*{$1$-$41$} $65.62$ $63.42$
@@ 5.2.3 Analysis
Self $51.23$ $65.24$ $43.31$ $51.05$ $65.77$ $43.04$ $51.14$ $65.50$ $43.17$
Synthesized $57.11$ $66.31$ $52.16$ $56.25$ $64.35$ $50.79$ $56.68$ $64.23$ $51.44$
   - filter $56.78$ $65.16$ $53.26$ $56.08$ $64.09$ $50.81$ $56.43$ $64.62$ $52.03$
   vqa filter $56.44$ $63.95$ $51.11$ $55.89$ $64.07$ $50.41$ $56.16$ $64.01$ $50.76$
   filter local $56.75$ $63.87$ $51.61$ $56.34$ $64.06$ $51.22$ $56.55$ $63.97$ $51.41$
@@
Self $51.23$ $65.24$ $43.31$ $51.05$ $65.77$ $43.04$ $51.14$ $65.50$ $43.17$
   loop1 $51.91$ $63.48$ $45.02$ $52.35$ $64.98$ $45.03$ $52.13$ $64.23$ $45.03$
   loop2 $52.50$ $63.43$ $45.66$ $52.43$ $63.81$ $45.70$ $52.47$ $63.62$ $45.68$
   loop3 $52.89$ $62.45$ $46.86$ $52.78$ $63.00$ $46.52$ $52.84$ $62.73$ $46.69$
   loop4 $54.02$ $62.24$ $48.45$ $54.37$ $62.89$ $48.78$ $54.20$ $62.56$ $48.62$